If the coefficient of friction is 0.3 and Contact angle is 850, determine the transmitted


breaking torque for the following shoe brake.

1
Expert's answer
2022-01-05T02:27:01-0500

breaking torque "T_b = F_t \times r"


 "T_b = \mu N \times r"


"T_b = 0.3 N \times r"


Breaking Torque:

Brake torque is essentially the power of the braking system. The brake caliper acts on the disc at a certain distance from the hub center, known as the effective radius. The force exerted by the caliper, multiplied by the effective radius of the system equals the brake torque.